Is the sum or product of two irrational numbers always irrational?


1
Expert's answer
2021-01-19T18:10:38-0500

No, it isn't. Here is the example:


"\sqrt{2}" is irrational number.

"-\sqrt{2}" is irrational number.

And "1/\sqrt{2}" is irrational number too.


But:

"\sqrt{2}+(-\sqrt{2})=0" -- rational number.

"\sqrt{2}*(1/\sqrt{2})=1" -- rational number too.